OpenStudy (anonymous):

find the critical points of f(x y) = 4+x^3+y^3 -3xy

OpenStudy (wikiemol):

the critical points are where the differential of f(x y) = 0. so you want to know where the partial of f with respect to x and the partial of f with respect to y are 0. In other words you want to solve the system of equations: \[0 =3x^2 - 3y\] \[0 = 3y^2 - 3x\]

OpenStudy (anonymous):

\[x(1-x ^{-3/4})=0 and y(1-y ^{-3/4})\]

OpenStudy (anonymous):

so x=0 or x=1 and y=0 or y=1 then what

OpenStudy (wikiemol):

you see which values correspond to which by plugging either x or y into one of the equations and the points you get are the answers.

OpenStudy (wikiemol):

sorry, then you have to plug those points into the original function f(x,y) and THAT is your answer.

OpenStudy (wikiemol):

so in this case you would plug in (0,0) and (1,1) and you end up with the critical points (0,0,4) and (1,1,1)

OpenStudy (anonymous):

so why wouldn't we use (0,1) or (1,0)

OpenStudy (wikiemol):

well you would plug say x = 0 into one of the equations in the partial derivatives. so in the equation \[0 = 3x^2 - 3y\] if x =0 then y = 0

OpenStudy (anonymous):

oh thanks wish i could give you too medals

OpenStudy (wikiemol):

haha thanks. and no problem.